OpenStudy (anonymous):

factoring when k is not equal to one... 7k^2-26k+15..... show how to get answer with steps? thank you!

OpenStudy (domebotnos):

7k^2 - 26k + 15 Find what factors equal what 7*15 = 105 when added. -5 and -21 work. (k - 5)(k - 21) ___ ___ 7 7 7 goes into -21 evenly but not -5. So: (7k - 5)(k - 3) You can solve to see if it is right. (7k - 5)(k - 3) 7k^2 - 21k - 5k + 15 7k^2 - 26k + 15 The answer is (7k - 5)(k - 3).

OpenStudy (domebotnos):

Do u understand that?

OpenStudy (anonymous):

yes thank you SO much!!!

OpenStudy (anonymous):

I have another Im stuck on if its okay with you

OpenStudy (domebotnos):

Is it factoring again?

OpenStudy (anonymous):

yes

OpenStudy (domebotnos):

Go ahead :)

OpenStudy (anonymous):

8x^2-44x+60

OpenStudy (domebotnos):

8x^2 - 44x + 60 8*60=480 -20*-24 = 480 -20-24 = -44 (x - 20)(x - 24) _____ ______ 8 8 (8x - 20)(x - 3) Check: (8x - 20)(x - 3) 8x^2 - 24x - 20x + 60 8x^2 - 44x + 60 The answer is (8x - 20)(x - 3)

OpenStudy (domebotnos):

Do u get it?

OpenStudy (anonymous):

Oh okay! I get it! Thanks for the step by step! My teacher doesnt do that haha

OpenStudy (anonymous):

THANK YOU! <3

OpenStudy (domebotnos):

Np. :)

OpenStudy (domebotnos):

Medal?